A ladder is leaning against a building so that the distance from the ground to the top of the ladder is 1 foot less than the length of the ladder. Find the length of the ladder if the distance from the bottom of the ladder to the building is 7 ft.
Answer by stanbon(75887)

A ladder is leaning against a building so that the distance from the ground to the top of the ladder is 1 foot less than the length of the ladder. Find the length of the ladder if the distance from the bottom of the ladder to the building is 7 ft.
--------------
Sketch the picture.
You have a right triangle with
height = h
hypotenuse = ladder = h+1
base = 7 ft
----------------------
Use Pythagoras to solve for "h"::
(h+1)^2 = h^2 + 7^2
------
h^2 + 2h + 1 - h^2 = 49
2h + 1 = 49
2h = 48
height = 24 ft
-----------
Cheers,
Stan H.
-------------